OpenStudy (anonymous):

Q.8) 1.294 times 10 to the second power equals 129.4 A. True B. False

Parth (parthkohli):

What do you think?

Parth (parthkohli):

\[1.294 \times 10^2 \Longrightarrow 1.294 \times 100 \]

OpenStudy (anonymous):

dont know

Parth (parthkohli):

What is \(1.294 \times 100\)?

OpenStudy (anonymous):

1294

OpenStudy (anonymous):

when you multiply a decimal number by 10 a digit from right of decimal point shifts towards left

Parth (parthkohli):

No no

OpenStudy (anonymous):

its false

Parth (parthkohli):

No.

Parth (parthkohli):

\[1.294 \times 100 \implies 1.294 \times 10 \times 10 \implies 12.94 \times 10 \implies ? \]

OpenStudy (anonymous):

so its true

Parth (parthkohli):

Yeah.